OpenStudy (kmullis6):

Please help me! Will fan & give medal! Directions: Write an equation for the problem and then solve. Change this phrase into symbols. Nine more than "twice the value of a number decreased by 6".

OpenStudy (dumptybaby):

Equation: 2X -3 X= 2/3

OpenStudy (kmullis6):

Can you explain for me please?

OpenStudy (dumptybaby):

Here you go

OpenStudy (tkhunny):

Dumpty cannot explain it, because it makes no sense. Order of Operations Version Nine more than "twice the value of a number decreased by 6". 1) A Number, x is popular 2) Twice the value of a number, 2x 3) twice the value decreased by 6, 2x - 6 4) Nine more than... 9 + 2x - 6 That's an expression. There is no equation and thus, there is nothing to solve.

OpenStudy (tkhunny):

Why did we decide it was "= 0"? That's not in the problem statement.

OpenStudy (kmullis6):

I'm not really sure. Is it because it can't be solved?

OpenStudy (dumptybaby):

Oh. Yeah you make sense then it has no answer.

OpenStudy (dumptybaby):

It's a statement you would need to have one amount known to know the other

OpenStudy (kmullis6):

Simplified it is 2x + 3 right? Or is that even relevant?

OpenStudy (dumptybaby):

It's 2x - 3 and yes of course it is. Thai equation is your final answer.

OpenStudy (kmullis6):

But how is it 2x "-" 3?

OpenStudy (tkhunny):

It isn't 9 - 6 = +3 Don't let a typo mess with your head. Work with confidence and FIX errors that you encounter.

OpenStudy (dumptybaby):

It's actually 2(X-6)-9 which gives 2X - 12+9 which gives 2X - 3

OpenStudy (dumptybaby):

Very sorry I meant 2(X-6) +9 **** in the original equation

OpenStudy (kmullis6):

@dumptybaby I'm sorry but I guess I'm not getting it cus wouldn't 9 be first? @tkhunny so what are you saying is the final answer?

OpenStudy (dumptybaby):

9 + 2 (X-6) gives 9 + 2x -12 which gives the final answer 2X - 3

OpenStudy (kmullis6):

Oh ok, thanks!!

OpenStudy (dumptybaby):

sorry for any confusion

OpenStudy (kmullis6):

It's fine!

OpenStudy (tkhunny):

No, it's not "actually 2(X-6)-9". English can be quite ambiguous. I followed strictly the rues of Order of Operations as taught in most standard algebra books. You can get the expression you suggest, but not by following those rules. If you would like to follow some other set of rules, please state what rules they are. "Whatever seems to make sense in your head" is not a very good rule.